OpenStudy (real.420):

Three bottles poured into bigger container. individual bottles have a mean of 15 ounces and standard deviation of .08. what is the standard deviation of the volume of liquid emptied to bigger container?

OpenStudy (kropot72):

The variance of the individual bottles is 0.08^2 = 0.0064. The variance of the volume of liquid in the bigger container is 0.0064 * 3 = 0.0192. The standard deviation of the volume of liquid in the bigger container is given by: \[\large S.D.=\sqrt{0.0192}=0.1386\]